OpenStudy (anonymous):

WILL GIVE METAL TO THE BEST ANSWER!!!!! Is the simplified form of 2square root of 3 - 2square root of 3 rational? Yes No

OpenStudy (anonymous):

@mtbender74

OpenStudy (anonymous):

\[2\sqrt{3}-2\sqrt{3}\] or \[2\sqrt{3-2\sqrt{3}}\]

OpenStudy (anonymous):

so it it rational or not ration @mtbender74

OpenStudy (anonymous):

rational*

OpenStudy (anonymous):

which of the two that i posted is the way the problem is written?

OpenStudy (anonymous):

sorry...it didn't take the first part of my above post...i was trying to ask which one the question was

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

if it is the first one, then \[2\sqrt{3} - 2\sqrt{3} = 0\] which is certainly rational...